# Centrality in Collaboration: A Novel Algorithm for Social Partitioning Gradients in Community Detection for Multiple Oncology Clinical Trial Enrollments

### v1

@@ 1 Introduction
When cancer patients complete standard treatments at Princess Margaret Cancer Centre and have not responded with being cured or in remission, they become candidates for clinical trials. These clinical trials are regulated studies registered by Health Canada [^1] For more information, see https://www.canada.ca/en/health-canada/services/clinical-trials.html as opposed to quality assurance studies [^2] Quality assurance studies in the context of medical studies are studies which look at drugs which are already approved for use, but the goals are focused on other aspects of care such as drug delivery or quality of care.. Patients who qualify may have been screen failures for other trials, have experienced progressive disease, or are receiving maintenance therapy and have been referred to a clinical trial by their oncologist or most responsible physician. Ground truth shows that collaboration networks between oncologists is a primary factor for further engagement in subsequent clinical trials by patients after completion of the given clinical trial that they are enrolled in. A possible approach to understanding the structure of these collaboration networks is through use of social network analysis (SNA) and community detection algorithms.
@@
Application of community detection algorithms with oncology clinical trial data has been preformed in the past. Georgiev et al2 (2011) applied the Girvan-Newman3 (2002) algorithm and noted a lack of cohesion among researchers who studied treatments for multiple myeloma. Haq and Wang4 (2016) applied the Louvain algorithm (by Blondel et al5 (2008)) to identify communities of cancer patients with significantly different survival curves. The present study applies SNA, and compares multiple community detection algorithms to identify collaboration networks between oncologists through the interventions studied in clinical trials via enrollment data of patients in multiple, nonconcurrent clinical trials at Princess Margaret Cancer Centre (PM) in Toronto, Canada. Inspired by work from Gorgiev et al (2011), Haq and Wang (2016), Ostovari and Yu6 (2019) and Bissoyi and Patra7 (2020) this research considers the Girvan-Newman and Louvain algorithms and compares them to an author-developed algorithm, referred to as “Smith-Pittman” [^3] Named after the author and his co-supervisor, Tyler Pittman., to identify collaboration networks between clinical trials classified by intervention.
@@ 2.1 The Data
The data is a result of multiple data source integration, which was undertaken from the PM Cancer Registry and Clinical Research Record. The data is anonymized, and spans patient enrollments in oncology clinical trials between January 1, 2016 and December 31, 2018. In this time period, there were 2970 patients enrolled in 515 clinical trials involving 41 principal investigators. For the identification of collaboration networks between oncologists, the analytic sample only consists of patients who were enrolled in more than one clinical trial within the time period studied. The resulting analytic sample consists of 389 patients enrolled in 288 clinical trials. Among these clinical trials, some interventions can be classified into broader categories of targeted therapies, or immunotherapy. This has been identified in the data with “T:” and “I:” prefixes respectively. The clinical trials were classified by intervention type, presenting as 16 distinct intervention types among 470 patient enrollments. With this classification, the patient referral graph is constructed (see Figure 2).

### v2

@@ 1 Introduction
When cancer patients complete standard treatments, and have not responded with being cured or in remission, they become candidates for clinical trials. These clinical trials are regulated studies registered by Health Canada [^1] For more information, see https://www.canada.ca/en/health-canada/services/clinical-trials.html as opposed to quality assurance studies [^2] Quality assurance studies in the context of medical studies are studies which look at drugs which are already approved for use, but the goals are focused on other aspects of care such as drug delivery or quality of care.. Patients who qualify may have been screen failures for other trials, have experienced progressive disease, or are receiving maintenance therapy and have been referred to a clinical trial by their oncologist or most responsible physician. Ground truth shows that collaboration networks between oncologists is a primary factor for further engagement in subsequent clinical trials by patients after completion of the given clinical trial that they are enrolled in. A possible approach to understanding the structure of these collaboration networks is through use of social network analysis (SNA) and community detection algorithms.
@@
Application of community detection algorithms with oncology clinical trial data has been preformed in the past. Georgiev et al2 (2011) applied the Girvan-Newman3 (2002) algorithm and noted a lack of cohesion among researchers who studied treatments for multiple myeloma. Haq and Wang4 (2016) applied the Louvain algorithm (by Blondel et al5 (2008)) to identify communities of cancer patients with significantly different survival curves. The present study applies SNA, and compares multiple community detection algorithms to identify collaboration networks between oncologists through the interventions studied in clinical trials via enrollment data of patients in multiple, nonconcurrent clinical trials. Inspired by work from Gorgiev et al (2011), Haq and Wang (2016), Ostovari and Yu6 (2019) and Bissoyi and Patra7 (2020) this research considers the Girvan-Newman and Louvain algorithms and compares them to an author-developed algorithm, referred to as ”Smith-Pittman” [^3] Named after the author and his co-supervisor, Tyler Pittman., to identify collaboration networks between clinical trials classified by intervention.
@@ 2.1 The Data
The data is simulated oncology clinical trials. There were 2970 patients enrolled in 515 clinical trials involving 41 principal investigators. For the identification of collaboration networks between oncologists, the analytic sample only consists of patients who were enrolled in more than one clinical trial within the time period studied. The resulting analytic sample consists of 389 patients enrolled in 288 clinical trials. Among these clinical trials, some interventions can be classified into broader categories of targeted therapies, or immunotherapy. This has been identified in the data with “T:” and “I:” prefixes respectively. The clinical trials were classified by intervention type, presenting as 16 distinct intervention types among 470 patient enrollments. With this classification, the patient referral graph is constructed (see Figure 2).